OpenStudy (anonymous):

Determine the mean root square speed of CO2 molecules that have the average kinetic energy of 4.21 x 10^-21 in m/s

OpenStudy (anonymous):

@ParthKohli can you help?

Parth (parthkohli):

Oh, never mind.

Parth (parthkohli):

For one CO2 molecule, the kinetic energy is \(\frac{1}{2} \times m \times v^2.\) Do you know the mass of a CO2 molecule?

OpenStudy (anonymous):

yes 44.01

Parth (parthkohli):

Just take that as 44.

OpenStudy (anonymous):

okay

Parth (parthkohli):

But you need to convert it to kilograms. Can you do that?

OpenStudy (anonymous):

.044?

Parth (parthkohli):

OK, nope... First of all, do you know how you would convert from atomic mass units to grams?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

will it be 44 times avagadros number?

OpenStudy (anonymous):

to change it into kg

Parth (parthkohli):

Alright, look: 44 g = Avogadro's No. * 44 u So 44 u = (44/Avogadro's No.) grams.

OpenStudy (anonymous):

I got 7.3

Parth (parthkohli):

* 10^(-25).

OpenStudy (anonymous):

well 7.306 x 10^-23

Parth (parthkohli):

(44/6*10^23*10^3) kg = 44/(6*10^26) kg = 7.3 * 10^(-26) kg I've been making so many errors. Ugh. Anyway...

Parth (parthkohli):

\[v^2 = \frac{\rm KE}{\frac{1}{2}m}\]

OpenStudy (anonymous):

so that answer is the m in the equation?

Parth (parthkohli):

No, it is \(v\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

v^2 = 2E/m = 2 (4.21x10^-21 J)/7.31 x 10^-26 kg = 57600 m^2/s^2 v = 240 m/s is that correct?

OpenStudy (anonymous):

v^2 = 2E/m = 2 (4.21x10^-21 J)/7.31 x 10^-26 kg = 57600 m^2/s^2 v = 240 m/s is that correct?

OpenStudy (anonymous):

v^2 = 2E/m = 2 (4.21x10^-21 J)/7.31 x 10^-26 kg = 57600 m^2/s^2 v = 240 m/s is that correct?

OpenStudy (anonymous):

v^2 = 2E/m = 2 (4.21x10^-21 J)/7.31 x 10^-26 kg = 57600 m^2/s^2 v = 240 m/s is that correct?